# Auto-Regressively Generating Multi-View Consistent Images

### v1

@@ Abstract
Generating multi-view images from human instructions is crucial for 3D content creation. The primary challenges involve maintaining consistency across multiple views and effectively synthesizing shapes and textures under diverse conditions. In this paper, we propose the Multi-View Auto-Regressive (MV-AR) method, which leverages an auto-regressive model to progressively generate consistent multi-view images from arbitrary prompts. Firstly, the next-token-prediction capability of the AR model significantly enhances its effectiveness in facilitating progressive multi-view synthesis. When generating widely-separated views, MV-AR can utilize all its preceding views to extract effective reference information. Subsequently, we propose a unified model that accommodates various prompts via architecture designing and training strategies. To address multiple conditions, we introduce condition injection modules for text, camera pose, image, and shape. To manage multi-modal conditions simultaneously, a progressive training strategy is employed. This strategy initially adopts the text-to-multi-view (t2mv) model as a baseline to enhance the development of a comprehensive X-to-multi-view (X2mv) model through the randomly dropping and combining conditions. Finally, to alleviate the overfitting problem caused by limited high-quality data, we propose the “Shuffle View” data augmentation technique, thus significantly expanding the training data by several magnitudes. Experiments demonstrate the performance and versatility of our MV-AR, which consistently generates consistent multi-view images across a range of conditions and performs on par with leading diffusion-based multi-view image generation models.
@@ 1 Introduction
However, new framework introduces new practical issues. This study is dedicated to the examination of the issues  [^1] This study focuses on addressing the AR model’s issues associated with generating multi-view images. The resolution of fundamental issues inherent to AR models, including their unidirectional nature and discrete encoding, is designated for future work. associated with the application of AR models in the context of multi-view image generation and proposes solutions thereto. We delineate three main issues that inhibit the effective development of AR models within multi-view image generation: insufficient conditions, limited high-quality data, and cumulative error. We systematically address these problems from two distinct perspectives: architectures and training. We devise effective methods for conditional injection that apply to text [36, 32], camera pose [15], reference image [21, 22], and shape [32, 44] in our MV-AR. Specifically, we practically demonstrate that the in-context image condition is not suitable for the current AR model. We design a precise Image Warp Controller (IWC) module that extracts features of the overlap between the current view and the previous views. IWC inserts condition token-by-token, thereby ensuring accurate image control. Next, we progressively drop or combine multi-modal conditions, so that the model can handle multi-modal conditions simultaneously after training. Finally, to address the challenge of limited training data, we formulate a data augmentation termed “Shuffle Views” and a progressive learning strategy to enrich the training data and mitigate the risk of overfitting. Utilizing the aforementioned solutions, we formulate a series of guidelines for the AR Model in multi-view image generation and establish a new baseline for its application.
@@ 3.3 Training
Progressive Learning. Utilizing the aforementioned loss function Eq. 8 and the “ShufV” data augmentation Eq. 9, we can successfully train a text-to-multi-view (t2mv) model. This t2mv model employs SSA to analyze text input and efficiently produce multi-view images that correspond to the given text. We use the t2mv model as a baseline to train the X-to-multi-view (X2mv) model to facilitate the versatility of the MV-AR in alternative conditions, such as images and shapes.
During training of the X2mv model, the text condition is randomly dropped, while other conditions are randomly combined, as shown in Figure 2. When the text prompt is dropped, it is replaced by a statement that does not pertain to the target image. For example, a command such as “Generate multi-view images of the following $<$img$>$” may be utilized. In this context, “$<$img$>$” signifies that a reference image will be combined after the text. In cases where the subsequent element is a geometric shape, “$<$img$>$” is substituted with “$<$shape$>$”. The probability of a condition dropping and combining increases linearly from 0 to 0.5 as the number of training iterations increases, and remains 0.5 in subsequent training. This escalation is confined to the initial 10k training iterations. This progressive learning allows the model to be influenced by new conditions introduced during training while maintaining a certain level of adherence to the text prompts.
@@ 5 Conclusion
In this study, we introduce an approach for auto-regressively generating multiview images. The motivation is to ensure that, during the generation of the current view, the model can extract efficient guidance information from all preceding views, thus enhancing the multi-view consistency. Based on this idea, we develop the MV-AR model, which progressively produces multi-view images based on multi-modal conditions. Specifically, we initially design several conditional injection modules for four conditions: text, image, camera pose, and geometry. Subsequently, we implement progressive learning to enable the MV-AR to manage these conditions concurrently. Finally, we employ “Shuffle View” data augmentation to expand the dataset, mitigating the inherent overfitting issues associated with the AR model when applied to limited training data. As a result of these novel designing, our MV-AR model achieves generation performance that is comparable to the state-of-the-art diffusion-based multi-view image generation and establishes itself as the first multi-view image generation model capable of handling multi-modal conditions.

### v2

@@ Abstract
Generating multi-view images from human instructions is crucial for 3D content creation. The primary challenges involve maintaining consistency across multiple views and effectively synthesizing shapes and textures under diverse conditions. In this paper, we propose the Multi-View Auto-Regressive (MV-AR) method, which leverages an auto-regressive model to progressively generate consistent multi-view images from arbitrary prompts. Firstly, the next-token-prediction capability of the AR model significantly enhances its effectiveness in facilitating progressive multi-view synthesis. When generating widely-separated views, MV-AR can utilize all its preceding views to extract effective reference information. Subsequently, we propose a unified model that accommodates various prompts via architecture designing and training strategies. To address multiple conditions, we introduce condition injection modules for text, camera pose, image, and shape. To manage multi-modal conditions simultaneously, a progressive training strategy is employed. This strategy initially adopts the text-to-multi-view (t2mv) model as a baseline to enhance the development of a comprehensive X-to-multi-view (X2mv) model through the randomly dropping and combining conditions. Finally, to alleviate the overfitting problem caused by limited high-quality data, we propose the “Shuffle View" data augmentation technique, thus significantly expanding the training data by several magnitudes. Experiments demonstrate the performance and versatility of our MV-AR, which consistently generates consistent multi-view images across a range of conditions and performs on par with leading diffusion-based multi-view image generation models. The code and models are released here.
@@ 1 Introduction
However, new framework introduces new practical issues. This study is dedicated to the examination of the issues  [^1] This study focuses on addressing the AR model’s issues associated with generating multi-view images. The resolution of fundamental issues inherent to AR models, including their unidirectional nature and discrete encoding, is designated for future work. associated with the application of AR models in the context of multi-view image generation and proposes solutions thereto. We delineate three main issues that inhibit the effective development of AR models within multi-view image generation: insufficient conditions, limited high-quality data, and cumulative error. We systematically address these problems from two distinct perspectives: architectures and training. We devise effective methods for conditional injection that apply to text [36, 32], camera pose [15], reference image [21, 22], and shape [32, 44] in our MV-AR. Specifically, we practically demonstrate that the in-context image condition is not suitable for the current AR model. We design a precise Image Warp Controller (IWC) module that extracts features of the overlap between the current view and the previous views. IWC inserts condition token-by-token, thereby ensuring accurate image control. Next, we progressively drop or combine multi-modal conditions, so that the model can handle multi-modal conditions simultaneously after training. Finally, to address the challenge of limited training data, we formulate a data augmentation termed “Shuffle Views" and a progressive learning strategy to enrich the training data and mitigate the risk of overfitting. Utilizing the aforementioned solutions, we formulate a series of guidelines for the AR Model in multi-view image generation and establish a new baseline for its application.
@@ 3.3 Training
Progressive Learning. Utilizing the aforementioned loss function Eq. 8 and the “ShufV" data augmentation Eq. 9, we can successfully train a text-to-multi-view (t2mv) model. This t2mv model employs SSA to analyze text input and efficiently produce multi-view images that correspond to the given text. We use the t2mv model as a baseline to train the X-to-multi-view (X2mv) model to facilitate the versatility of the MV-AR in alternative conditions, such as images and shapes.
During training of the X2mv model, the text condition is randomly dropped, while other conditions are randomly combined, as shown in Figure 2. When the text prompt is dropped, it is replaced by a statement that does not pertain to the target image. For example, a command such as “Generate multi-view images of the following $<$img$>$" may be utilized. In this context, “$<$img$>$" signifies that a reference image will be combined after the text. In cases where the subsequent element is a geometric shape, “$<$img$>$" is substituted with “$<$shape$>$". The probability of a condition dropping and combining increases linearly from 0 to 0.5 as the number of training iterations increases, and remains 0.5 in subsequent training. This escalation is confined to the initial 10k training iterations. This progressive learning allows the model to be influenced by new conditions introduced during training while maintaining a certain level of adherence to the text prompts.
@@ 5 Conclusion
In this study, we introduce an approach for auto-regressively generating multiview images. The motivation is to ensure that, during the generation of the current view, the model can extract efficient guidance information from all preceding views, thus enhancing the multi-view consistency. Based on this idea, we develop the MV-AR model, which progressively produces multi-view images based on multi-modal conditions. Specifically, we initially design several conditional injection modules for four conditions: text, image, camera pose, and geometry. Subsequently, we implement progressive learning to enable the MV-AR to manage these conditions concurrently. Finally, we employ “Shuffle View" data augmentation to expand the dataset, mitigating the inherent overfitting issues associated with the AR model when applied to limited training data. As a result of these novel designing, our MV-AR model achieves generation performance that is comparable to the state-of-the-art diffusion-based multi-view image generation and establishes itself as the first multi-view image generation model capable of handling multi-modal conditions.